A grating has 10,000 lines per cm. What is the maximum number of principal  maxima that can be formed for light of wavelength 5×10-5 cm?


1
Expert's answer
2021-06-07T09:41:13-0400

Solution. The distance between slits is 



"d=\frac {1cm}{10000}=10^{-4}cm=10^{-6}m"

Once a value for the diffraction grating’s slit spacing d has been determined, the angles for the sharp lines can be found using the equation



"d sin\theta =m\lambda"

where Θ is angle between normal and maximum direction; m is number of principal maxima; λ=487nm is wavelength. Therefore



"m=\frac{d sin\theta}{\lambda}"

Using the properties of sine function



"sin\theta \eqslantless 1"

Hense



"m\eqslantless\frac{d }{\lambda}""\frac{d }{\lambda} =\frac{10^{-6}}{5\times 10^{-7}}=2"

As result get the maximum number of principal maxima is equal to 2.


Answer. 2.